Hi and thank you in advanced for your help. I'm in a pre-algebra class and I missed class yesterday and I'm supposed to finish 2 work sheets about functions. My pre-algebra teacher gave me the definition but I don't fully understand it. Could you please explain it to me in a way that I can understand? Also in one of the work sheets it says to "find four solutions to each equation" and then it gives me this:
1. y = x - 5
What am I supposed to do with it?

Answer by Edwin McCravy(8908)

Hi and thank you in advanced for your help. I'm in a pre-algebra class and I missed class yesterday and I'm supposed to finish 2 work sheets about functions. My pre-algebra teacher gave me the definition but I don't fully understand it. Could you please explain it to me in a way that I can understand? Also in one of the work sheets it says to "find four solutions to each equation" and then it gives me this:
1. y = x - 5
What am I supposed to do with it?

At random just pick any number you like for one of the letters,
plug it into the equation, and solve for the other letter. 
Then write the ordered pair with the coordinate you picked and
the one youi got.

For instance, suppose I pick the number 6 for x, Then I substitute
6 for x, like this
   
   y = x - 5
   y = 6 - 5
   y = 1

So I put the number I picked first, since x comes first,
in an ordered pair, and the number I got for y second,
since y comed after x.  So one solution is (6, 1).

suppose I pick the number -3 for y, Then I substitute
-3 for y, like this
   
   y = x - 5
  -3 = x - 5
   2 = x

So I put the number I GOT first, since x comes first,
in an ordered pair, and the number I picked for y second,
since y comed after x.  So one solution is (2, -3).
 
You can find as many solutions as you want by this method.
It doesn't matter what number you pick, and it doesn't
matter which letter you substitute it for.

Edwin